OpenStudy (anonymous):

A cement-mixing truck holds 9 cubic yards of cement. The truck delivers all of the cement in equal batches to 2 different locations. How many cubic yards of cement did the truck deliver to each location? Choose all answers that are correct.

OpenStudy (anonymous):



OpenStudy (anonymous):

@k_lynn

OpenStudy (k_lynn):

Since there are two equal sections that are being passed out, you need to divide 9 by 2. \[9\div2\]Then you'll convert the decimal to a fraction.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4.5

OpenStudy (k_lynn):

Right. and what's 4.5 as a fraction?

OpenStudy (anonymous):

4/5 or 5/4

OpenStudy (k_lynn):

Not quite. Since the 5 in the decimal number is in the tenths place, you create a fraction that looks like this. \[4\frac{ 5 }{ 10 }\]Then you reduce.

OpenStudy (anonymous):

hm...

OpenStudy (k_lynn):

This site is really helpful for me when I don't understand something. This is a video that tells you how to create decimals into fractions. https://www.khanacademy.org/math/cc-fourth-grade-math/cc-4th-fractions-topic/cc-4th-decimals-fractions/v/converting-decimals-to-fractions-1-ex-1

OpenStudy (anonymous):

cool thx

OpenStudy (k_lynn):

You're welcome. Tell me if you understand it or not when you're done. I'll see if I can find something that will explain more, if it doesn't seem to work.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

\[\frac{ 9 }{ 2 } =5\]

OpenStudy (k_lynn):

You're really close. But if you count up by twos 10 times like this: 2,4,6,8,10 There's no 9 in there. And there's no odd numbers. So \[9\div2\]has to be a decimal.

OpenStudy (k_lynn):

What number can you multiply by 2 to get 9?

OpenStudy (k_lynn):

http://www.mathsisfun.com/converting-decimals-fractions.html This site also might be helpful.

OpenStudy (anonymous):

sorry me cat got lost so ya had look for him but he came to me and all he wanted was to race me to his food bowl but anyway thx

OpenStudy (k_lynn):

lol that's ok.

OpenStudy (anonymous):

lol name is ender

OpenStudy (anonymous):

btw there is 2 answers in the qustion

OpenStudy (k_lynn):

okie doke. Thanks.

OpenStudy (anonymous):

yah and I have about 3-2 more Q.after sorry

OpenStudy (anonymous):

2 * 4.5 = 9

OpenStudy (k_lynn):

Right :). Now since 4.5 is your decimal, you can make a fraction. the 5 in 4.5 is in the tenths place. So you put the 5 over 10 to make a fraction. \[4\frac{ 5 }{ 10}\]But you can reduce that. How many times does 5 go into 10?

OpenStudy (anonymous):

brb

OpenStudy (k_lynn):

okie doke.

OpenStudy (anonymous):

back sorry

OpenStudy (anonymous):

2 times

OpenStudy (k_lynn):

Right, so now the fraction is \[4\frac{ 1 }{ 2 }\]

OpenStudy (k_lynn):

right?

OpenStudy (anonymous):

yes... i think

OpenStudy (k_lynn):

Yep, you got it. So C is an answer.

OpenStudy (anonymous):

brb me pc restart

OpenStudy (anonymous):

tag me

OpenStudy (k_lynn):

ok. @djspeed

OpenStudy (anonymous):

@k_lynn